Daily Events on 2015/04/28 (Tue)

last updated at  2015/04/27 13:28:41 JST


<< prev day

@

next day >>


Station

Pass No.
20150428-

Contact time (UT)

Max. Elev.
 (deg)

Events

AOS

LOS

USC20

0100

07:48:06

08:00:45

37.9

OP/OG upload
XRT: TI-CMD
EIS: TI-CMD
SOT: TI-CMD
DR playback (by OP)

USC20

0200

09:26:15

09:36:58

15.4

OP/OG upload, XRT:TI-CMD, SOT:TI-CMD, EIS:TI-CMD (backup)
AOCS: Upload pointing table
AOCS: Upload orbital element
EIS: Table Upload
DR playback (by OP)

SNT1

4010

10:15:14

10:28:37

39.9

SOT: Table Upload
XRT: Table Upload

Flare list(LM) : http://www.lmsal.com/solarsoft/last_events/


DR information

OP main: 2015/04/28 10:39:00 - 2015/04/30 10:13:00 (UT)

Remaining amount
(prediction)

3483 Mbits @ OP_START ( 2015/04/28 10:39:00 )


DR usage: The DR would become empty on Apr 29 at 3 UT if all telescopes took no images until that time.


SOT

XRT

EIS

DR allocation [%]

70

15

15

DR allocation [Mbits]

6065

1299

1299

DR TLMLMT [Mbits]

2259 @OP_START
-> 9018

1905

1905


daily note

SOT:
- IHOP254 Prominence Obs. (01-03 UT on 29th)
- Granulation Obs. (3hrs., 01-04 UT on 30th)
- AR Obs. @AR12333
- filament Obs.

XRT:
- support IHOP254
- AR obs. @12333
- filament Obs.

EIS:
- Sensitivity Monitoring (4hrs.) with Granulation Obs. (01-05 UT on 30th)
- support IHOP254
- AR Obs. @AR12333
- filament Obs. (11-17 UT on 28th)


CHIEF
TOHBAN

SUB
TOHBAN

CP

SOT CO

XRT CO

EIS CO

EVENING

Yasuda

Bamba

Bamba

Shimojo

Ishii/Lee

MORNING


operation files

pointing file : cmdpln/20150428/re-point_201504280100.txt